For a right triangle that has legs a and b and hypotenuse c, find the length of the side not given here. a = 5 cm b=12cm

Answer by Cromlix(4381)

This is a Pythagorean question.
The square on the hypotenuse is equal
to the the sum of the squares on the other two sides.
Therefore c^2 = a^2 + b^2
c^2 = 5^2 + 12^2
c^2 = 25 + 144
c^2 = 169
c = the square root of 169
c = 13 cm